# Unsupervised Text Embedding Space Generation Using Generative Adversarial Networks for Text Synthesis

### v1

@@ Abstract
Generative Adversarial Networks (GAN) is a model for data synthesis, which creates plausible data through the competition of generator and discriminator. Although GAN application to image synthesis is extensively studied, it has inherent limitations to natural language generation. Because natural language is composed of discrete tokens, a generator has difficulty updating its gradient through backpropagation; therefore, most text-GAN studies generate sentences starting with a random token based on a reward system. Thus, the generators of previous studies are pre-trained in an autoregressive way before adversarial training, causing data memorization that synthesized sentences reproduce the training data. In this paper, we synthesize sentences using a framework similar to the original GAN. More specifically, we propose Text Embedding Space Generative Adversarial Networks (TESGAN) which generate continuous text embedding spaces instead of discrete tokens to solve the gradient backpropagation problem. Furthermore, TESGAN conducts unsupervised learning which does not directly refer to the text of the training data to overcome the data memorization issue. By adopting this novel method, TESGAN can synthesize new sentences, showing the potential of unsupervised learning for text synthesis. We expect to see extended research combining Large Language Models with a new perspective of viewing text as an continuous space.
Keywords:
Text Embedding Space Text Synthesis Generative Adversarial Network Adversarial Learning Data Memorization Backpropagation.
@@ 1 Introduction
Generative Adversarial Networks (GAN) proposed by Goodfellow et al. (2014) [7] is a popular model for synthesizing data. GAN aims to generate plausible data, capturing real data distribution in an unsupervised way through the competition of generator and discriminator. When GAN was first proposed, it mainly focused on image synthesis, and extensive research was dedicated to achieving quality synthetic data results [2, 22, 13, 12]. The generator of GAN learns implicit density based on the loss of the discriminator without directly referencing to the training data; therefore, GAN can avoid data memorization that the model reproduces the training data. Also, GAN can synthesize various data with random noise instead of using specific starting point such as a start token.
@@
Applications of GAN to image synthesis have been widely studied, but it has had limitations to natural language generation. As natural language is composed of discrete tokens, the generator has an issue with updating its gradient directly through backpropagation. The backpropagation problem of text-based GAN was first discussed by Yu et al. (2017) [29] and a lot of follow-up text-GAN research tried to deal with the backpropagation problem, using gradient policy-based reinforcement learning with a reward system. Also, the previous text-GAN approaches have to pre-train the generator with supervised learning (autoregressive) before adversarial training due to the convergence issue of the generator [29]. Accordingly, we found that the generators of previous text-GAN approaches reproduce the training data (data memorization) when synthesizing text due to the pre-training and this is a major issue of the generative models.
@@
This paper suggests a novel framework, Text Embedding Space Generative Adversarial Networks (TESGAN), to enable backpropagation and avoid data memorization. The gradient backpropagation is possible in TESGAN because the model generates continuous text embedding spaces from random noise instead of discrete tokens. In addition, TESGAN does not require the supervised pre-trained generator adopted in previous research for text synthesis. In other words, TESGAN does not directly refer to the training data because it trains in an unsupervised way similar to the original GAN framework. Thus, data memorization does not arise while synthesizing sentences. We compute Fréchet BERT Distance (FBD) and Multi-sets-Jaccard (MSJ) proposed by Alihosseini et al. (2019) [1], which evaluate both quality and diversity of synthetic text simultaneously. Furthermore, TESGAN-based models show the best Language Model score (LM) [20, 4] compared to baselines. We also calculate data memorization ratio and show the synthesized sentences to assess the potential of unsupervised learning and continuous embedding spaces for text synthesis.
@@ 2 Related Works
The most common method of sentence generation is to use an autoregressive-based language model via teacher forcing [28]. For example, there are extensive studies on models using recurrent neural networks (RNN) with Long Short-Term Memory (LSTM) cells proposed by Hochreiter and Schmidhuber (1997) [11]. Using LSTM, Graves (2013) [8] successfully generated handwriting by predicting a sequence, and Wen et al. [27] synthesized sentences under specific conditions. Bowman et al. (2016) [3] generated text after learning text embedding spaces with an autoregressive-based LSTM model and variational autoencoder (VAE) architecture [15]. Policy Gradient with BLEU (PG-BLEU) computes the BLEU [21] of synthesized sentences and takes them as a reward when updating the generator with policy gradient.
@@
Many investigations were conducted to use GAN for text synthesis. Sequence GAN (SeqGAN) proposed by Yu et al. (2017) [29] attempted to solve the backpropagation problem by using gradient policy-based reinforcement learning with a reward system. However, since SeqGAN has reward sparsity issue, Lin et al. (2017) [18] proposed RankGAN which replaces the previous regression-based discriminator with a novel ranker. RankGAN trains the discriminator which gives a higher score to more realistic sentences. In addition, Che et al. (2017) [5] proposed Maximum Likelihood Augmented Discrete GAN (MaliGAN) that synthesizes text by minimizing Kullback-Leibler divergence [16]. TextGAN [30] proposed kernel-based moment-matching which forces empirical distributions of real and synthetic text using LSTM and Convolutional Neural Networks (CNN) for the generator and the discriminator, respectively. LeakGAN [9] used a similar architecture to TextGAN, but they alleviated sparseness and lack of intermediate information by providing leaked information from the discriminator. The aforementioned text-GAN models need the first token to synthesize text because of an autoregressive generator. TESGAN is distinct from these models in that it generates text embedding space directly from random noise without selecting tokens. Our TESGAN is the first text-GAN model which learns without an autoregressive generator.
@@ 3 Text Embedding Space GAN
TESGAN aims to generate seeds required to synthesize plausible text. The generated seeds (fake seeds) from the generator and the real seeds from the real text are passed over to the discriminators to train the GAN framework.
Figure 1: Illustration of the seed interpretation model. After pre-training the seed interpretation model using multi-turn sentences, the model is used to synthesize text after freezing the parameters. Regardless of pre-training and text synthesis, all data pass through the real seed making process (sigmoid embedding) before entering the model as input. The $[PAD]$ tokens after the $[SEP]$ tokens are omitted in the left part for easy understanding.
@@ 3.1 Seed for Text Synthesis
We denote a text sequence as $S=w_{1},...,w_{T}$ ($T$ is the sequence length). An autoregressive-based language model calculates the probability of the text sequence $S$ as a product of conditional probabilities. If we assume that $S$ is a complete sentence, then their sequences $D=S_{1},...,S_{N}$ ($N$ is the dialogue length) can be seen as multi-turn sentences. Let $S_{1}=w_{1},...,w_{m}$ and $S_{2},...,S_{N}=w_{m+1},...,w_{M}$ denote the first sentence and the subsequent sentences, respectively ($M$ is the total length of the multi-turn sentences $D$). The following sentences after the first sentence can be predicted from a product of conditional probabilities:
$\small p(S_{2},...,S_{N}|S_{1})=\prod_{i=m+1}^{M}p(w_{i}|w_{1},...,w_{i-1})$ (1)
In other words, the first sentence can generate subsequent sentences using an autoregressive-based language model trained with multi-turn sentences. Therefore, the first sentence can be used as a seed. Meanwhile, the generator of TESGAN generates the first sentence as a continuous embedding space instead of discrete tokens to enable gradient backpropagation. Consequently, a continuous embedding space of the first sentence is defined as a seed.
@@ 3.2 Seed Interpretation Model
The seed interpretation model $f_{\theta}(\cdot)$ is used to predict subsequent sentences a seed makes. Therefore, the seed interpretation model must be trained with multi-turn sentences in an autoregressive way as same as general language modeling, as shown in Figure 1. The model has been trained by data consisting of multi-turn sentences $D=S_{1},...,S_{N}$, and the maximum length of each sentence is the same as $L$, that is, the number of tokens in $D$ is $N*L$. When constructing the multi-turn sentence data, the $[CLS]$ special token is inserted only at the beginning of the first sentence, and each sentence is distinguished by adding the $[SEP]$ special token at the end. If the length of a tokenized sentence is less than $L$, the $[PAD]$ tokens are used for padding the sentence:
$\small\begin{gathered}S_{1}=w_{1}^{1},...,w_{|S_{1}|}^{1},...,w_{L}^{1}\\ (w_{1}^{1}=[CLS],w_{|S_{1}|}^{1}=[SEP],w_{|l>S_{1}|}^{1}=[PAD])\end{gathered}$ (2)
$\small\begin{gathered}S_{i(i>1)}=w_{1}^{i},...,w_{|S_{i}|}^{i},...,w_{L}^{i}\\ (w_{|S_{i}|}^{i}=[SEP],w_{|l>S_{i}|}^{i}=[PAD])\end{gathered}$ (3)
@@
where the $S_{1}$ is seed sentence. Let $H_{real}$ and $H_{fake}$ denote the real and the fake seeds, respectively, from TESGAN. As shown in Figure 1, the real seed is an embedding space of a sentence obtained by applying the sum of the token embedding layer and the positional embedding to the sigmoid function. Its dimension is (sequence length * embedding dimension). Any sentence can be used as a seed for text synthesis as long as the seed interpretation model is trained with multi-turn sentences because every sentence can exist before any other sentence. Therefore, all sentences can be used as a seed:
@@
$\small\begin{gathered}H_{real}=\sigma\big{(}W_{emb}(S_{1})+W_{pos}(S_{1})\big{)}\\ \approx\sigma\big{(}W_{emb}(S_{n})+W_{pos}(S_{n})\big{)}\end{gathered}$ (4)
$H_{real}$ from the real text can be seen as continuous spaces similar to an image. Therefore, if the generator can create $H_{fake}$ skillfully imitating distributions of $H_{real}$, $H_{fake}$ also can generate appropriate subsequent sentences (a.k.a synthetic text). No matter how excellent $H_{fake}$ are generated by the generator, it is useless if it cannot be interpreted; therefore, training the seed interpretation model is crucial. In other words, the seed interpretation model has to predict the next sentence $S_{n+1}$ properly in multi-turn cases as Equation 5. Consequently, text synthesis is conducted as the seed passes through the seed interpretation model to predict the next sentence as shown in Figure 1.
$\small S_{n+1}=f_{\theta}\Big{(}\sigma\big{(}W_{emb}(S_{n})+W_{pos}(S_{n})\big{)}\Big{)}$ (5)
GPT-2 [23] is used as the seed interpretation model and it is only used to synthesize sentences from seeds with frozen parameters. Thus the seed interpretation model never affects the training of the generator and the discriminator during adversarial training, and vice versa. More detailed specifications are explained in Appendix A.

### v4

@@ Abstract
Generative Adversarial Network (GAN) is a data synthesis model that creates plausible data through the competition between a generator and a discriminator. Although GAN has been extensively studied for image synthesis, it has inherent limitations when applied to natural language generation. This is because natural language is composed of discrete tokens, and the generator faces challenges in updating its gradient through backpropagation. Therefore, most text-GAN studies generate sentences starting with a random token (or prompt) based on a reward system. Thus, the generators of previous studies are pre-trained in an autoregressive manner before adversarial training, resulting in data memorization where synthesized sentences reproduce the training data. In this paper, we synthesize sentences using a framework similar to the original GAN. More specifically, we propose Text Embedding Space Generative Adversarial Networks (TESGAN), which generate continuous text embedding spaces instead of discrete tokens to address the gradient backpropagation problem. Furthermore, TESGAN conducts unsupervised learning that does not directly refer to the text of the training data to overcome the data memorization issue. Also, TESGAN enables unconditional text synthesis during the inference phase by using random noise instead of tokens or prompts for text synthesis. By adopting this novel method, TESGAN can synthesize new sentences, demonstrating the potential of unsupervised learning for text synthesis. We look forward to extended research that combines large-scale language models with a new perspective on viewing text as continuous spaces.
@@ 1 Introduction
Generative Adversarial Network (GAN), as proposed by Goodfellow et al. (2014), is a popular model for data synthesis. GAN is an unconditional data generation algorithm that aims to generate plausible data in an unsupervised manner by fostering competition between a generator and a discriminator to capture the real data distribution. When GAN was initially introduced, it primarily focused on image synthesis, and extensive research was conducted to achieve high-quality synthetic data results Arjovsky et al. (2017); Radford et al. (2015); Karras et al. (2018, 2021). Furthermore, GAN is commonly employed in the field of computer vision for data augmentation through image synthesis Sandfort et al. (2019); Bowles et al. (2018); Antoniou et al. (2018); Tran et al. (2021). The GAN generator learns implicit density based on the discriminator’s loss without direct reference to the training data. Consequently, GAN can prevent data memorization, where the model reproduces the training data. Additionally, GAN can synthesize various data by using random noise instead of a specific starting point, such as a designated start token.
@@
Similar to images, unconditional text generation can function as a data augmentation technique by generating new text that resembles a given dataset. It also has practical applications, such as creating new documents by generating fictitious text information suitable for direct use. Consequently, several studies have attempted to apply GAN to natural language, but they have encountered limitations in natural language generation. The challenge arises from the fact that natural language is composed of discrete tokens, making it challenging for the GAN generator to directly update gradients through backpropagation. The gradient backpropagation issue in text-based GANs was first discussed by Yu et al. (2017), and numerous subsequent text-GAN research efforts aimed to address this problem using gradient policy-based reinforcement learning with a reward system. Furthermore, the previous text-GAN approaches necessitated pre-training the generator with supervised learning (autoregressive) before adversarial training due to convergence issue with the generator Yu et al. (2017). Accordingly, we discovered that the generators of previous text-GAN approaches reproduce the training data (leading to data memorization) during text synthesis due to the autoregressive-based pre-training process, which becomes a significant issue in generative models.
@@
This paper introduces a novel framework known as Text Embedding Space Generative Adversarial Networks (TESGAN) [^1] https://github.com/ljm565/TESGAN, which enables backpropagation and prevents data memorization. TESGAN does not rely on a supervised, pre-trained autoregressive-based generator that generates discrete tokens for text synthesis. Our generator generates continuous text embedding spaces for text synthesis instead of discrete tokens, allowing training with gradient backpropagation. Furthermore, the fact that TESGAN deals with continuous spaces makes it possible for TESGAN’s generator to be trained within the original GAN framework to mimic the real text embedding space. Moreover, TESGAN enables unconditional text generation, as it does not require the selection of a starting token (or prompt) for text synthesis. Our seed interpretation model then synthesizes sentences by interpreting the imitated continuous text embedding space created by the generator. During sentence synthesis, data memorization does not occur because TESGAN does not directly refer to the training text data but only learns from the continuous text embedding space. We use two datasets to conduct performance evaluations and general applicability experiments based on synthetic text generated by TESGAN. To assess the quality and diversity of synthesized text, we employ evaluation metrics such as Fréchet BERT Distance (FBD), Multi-sets-Jaccard (MSJ) Alihosseini et al. (2019), Language Model score (LM) de Masson d’Autume et al. (2019); Caccia et al. (2020), and Self-BLEU (SBL) Zhu et al. (2018). In addition, we conducted human evaluations, and TESGAN achieved the highest average score. Lastly, we calculate the data memorization ratio and present the synthesized sentences to assess the potential of unsupervised learning and continuous embedding spaces for text synthesis.
@@ 2 Related Works
The most common method of text generation is to use an autoregressive-based language model via teacher forcing Williams and Zipser (1989). For example, extensive studies have been conducted on models using recurrent neural network (RNN) with Long Short-Term Memory (LSTM) cells Hochreiter and Schmidhuber (1997). Using LSTM, Graves (2013) successfully generated handwriting by predicting sequences, and Wen et al. (2015) synthesized sentences under specific conditions. Bowman et al. (2016) generated text after learning text embedding spaces with an autoregressive-based LSTM model and a variational autoencoder (VAE) architecture Kingma and Welling (2014). Policy Gradient with BLEU (PG-BLEU) calculates the BLEU Papineni et al. (2002) score of synthesized sentences and takes them as a reward when updating the generator using policy gradient.
@@
Numerous investigations have been conducted to utilize GANs for text synthesis. Sequence GAN (SeqGAN) Yu et al. (2017) attempted to address the backpropagation problem by employing gradient policy-based reinforcement learning with a reward system. However, SeqGAN faced a reward sparsity issue, leading Lin et al. (2017) to introduce RankGAN, which replaced the previous regression-based discriminator with a novel ranker. RankGAN trains the discriminator to assign higher scores to more realistic sentences. MaskGAN Fedus et al. (2018) utilized an LSTM-based generator to fill in masked parts of sentences with tokens during training. Since MaskGAN uses discrete tokens, gradient backpropagation is not possible for the generator. To overcome this challenge, the authors employed the actor-critic method, using the probabilities of candidate tokens from the discriminator as rewards during training. Che et al. (2017) proposed Maximum Likelihood Augmented Discrete GAN (MaliGAN), which synthesizes text by minimizing Kullback-Leibler divergence Kullback and Leibler (1951). LeakGAN Guo et al. (2018) alleviated issues related to sparseness and the lack of intermediate information by providing leaked information from the discriminator.
Several studies have aimed to address the gradient backpropagation problem without relying on reward-based reinforcement learning. TextGAN Zhang et al. (2017) introduced kernel-based moment-matching, which enforces empirical distributions of real and synthetic text by using LSTM and Convolutional Neural Networks (CNN) for the generator and the discriminator, respectively. Feature Mover GAN (FM-GAN) Chen et al. (2018) defined the feature-mover’s distance (FMD) and learned it by minimizing the FMD between real and fake sentences. Both TextGAN and FM-GAN utilized LSTM generators that generate discrete tokens using the soft-argmax trick instead of relying on reinforcement learning. Relational GAN (RelGAN) Nie et al. (2019) applied relational recurrent neural networks Santoro et al. (2018) and attempted to address the gradient backpropagation issue using Gumbel-softmax Jang et al. (2017). However, since these approaches employed autoregressive (e.g., LSTM) generators, they explicitly referenced the training text data during model training. Consequently, previous studies faced challenges in avoiding complete data memorization while synthesizing sentences due to an autoregressive generator. Lastly, Transformer-based Implicit Latent GAN (TILGAN) Diao et al. (2021) adopted a similar approach to TESGAN for addressing the gradient backpropagation issue based on the embedding space. However, TILGAN differs from TESGAN in that it was trained on a latent space compressed by the encoder, configured as an autoencoder transformer, and did not utilize embeddings learned from real language models.
Most of the aforementioned text-GAN models require the first token or prompt to synthesize text due to their autoregressive generators. TESGAN stands apart from these models as it generates text embedding space directly from random noise, eliminating the need for selecting tokens. Our TESGAN is the first text-GAN model that learns the real text embedding space without relying on an autoregressive generator.
@@ 3 Text Embedding Space GAN
TESGAN aims to generate the seeds required for synthesizing plausible text. These generated seeds (fake seeds) from the generator, along with the real seeds from the real text, are passed to the discriminators for training within the GAN framework. Once the training of TESGAN is complete, the pre-trained seed interpretation model synthesizes text using the fake seed created by the generator.
Figure 1: Illustration of the seed interpretation model. The seed interpretation model is pre-trained with multi-turn sentences before adversarial training (left). After pre-training, the model’s parameters are frozen, allowing it to synthesize text from the seed. The right figure implies that text can be synthesized from the seed. The $[PAD]$ tokens following the $[SEP]$ tokens are omitted in the left part for clarity.
@@ 3.1 Seed for Text Synthesis
We denote a text sequence as $S=w_{1},\ldots,w_{T}$ ($T$ is the sequence length). An autoregressive-based language model calculates the probability of the text sequence $S$ as a product of conditional probabilities. If we assume that $S$ is a complete sentence, then the sequence $D=S_{1},\ldots,S_{N}$ ($N$ is the dialogue length) can be viewed as multi-turn sentences. Let $S_{1}=w_{1},\ldots,w_{m}$ and $S_{2},\ldots,S_{N}=w_{m+1},\ldots,w_{M}$ denote the first sentence and the subsequent sentences, respectively ($M$ is the total length of the multi-turn sentences $D$). The subsequent sentences after the first sentence can be predicted from a product of conditional probabilities:
$\small p(S_{2},\ldots,S_{N}|S_{1})=\prod_{i=m+1}^{M}p(w_{i}|w_{1},\ldots,w_{i-1})$ (1)
In other words, the first sentence can generate subsequent text using an autoregressive-based language model trained with multi-turn sentences. Therefore, the first sentence can serve as a seed. Meanwhile, the generator of TESGAN generates the first sentence as a continuous embedding space instead of discrete tokens to enable gradient backpropagation. Consequently, a continuous embedding space of the first sentence is defined as a seed.
@@ 3.2 Seed Interpretation Model
We define a seed in Section 3.1, and the seed interpretation model $f_{\theta}(\cdot)$ is used to synthesize text based on the seed. To synthesize text, the seed interpretation model must first be trained with multi-turn sentences in an autoregressive manner, similar to general language modeling, before adversarial training, as shown in Figure 1 (left), with the following loss function:
$\mathcal{L}_{LM}=-\frac{1}{N}\sum^{N}_{n=1}{\log{\frac{exp(x_{n,y_{n}})}{\sum^{C}_{c=1}{exp(x_{n,c})}}}}$ (2)
This enables the generator to synthesize appropriate text by utilizing the fake embedding space it creates during the inference phase. More detailed explanations will be provided in the following section. As a result, the model has to be trained on data consisting of multi-turn sentences $D=S_{1},\ldots,S_{N}$, where each sentence has a maximum length of $L$, meaning the total number of tokens in $D$ is $N\times L$. When constructing the multi-turn sentence data, the special token $[CLS]$ is inserted only at the beginning of the first sentence, and each sentence is distinguished by adding the special token $[SEP]$ at the end. If the length of the tokenized sentence is less than $L$, the sentence is padded with the special token $[PAD]$:
$\small\begin{gathered}S_{1}=w_{1}^{1},\ldots,w_{|S_{1}|}^{1},\ldots,w_{L}^{1}\\ (w_{1}^{1}=[CLS],w_{|S_{1}|}^{1}=[SEP],w_{|l>S_{1}|}^{1}=[PAD])\end{gathered}$ (3)
$\small\begin{gathered}S_{i(i>1)}=w_{1}^{i},\ldots,w_{|S_{i}|}^{i},\ldots,w_{L}^{i}\\ (w_{|S_{i}|}^{i}=[SEP],w_{|l>S_{i}|}^{i}=[PAD])\end{gathered}$ (4)
@@
where $S_{1}$ and $S_{i}$ represent a seed sentence and subsequent text. Let $H_{real}$ denote the real seeds from TESGAN. As shown in Figure 1, the real seed is an embedding space of a sentence obtained by applying the sum of the token embedding and the positional embedding to the sigmoid function. Since most sentences can exist before others as long as the seed interpretation model is trained with multi-turn sentences, a significant portion of them can be used as seeds for text generation. Therefore, most of the sentences can be used as real seeds:
@@
$\small\begin{gathered}H_{real}=\sigma\big{(}W_{emb}(S_{1})+W_{pos}(S_{1})\big{)}\\ \approx\sigma\big{(}W_{emb}(S_{n})+W_{pos}(S_{n})\big{)}\in\mathbb{R}^{L\times d}\end{gathered}$ (5)
where $L$ and $d$ represent sequence length and embedding dimensions. $H_{real}$ from the real text can be viewed as continuous spaces, similar to images, and the well-pretrained seed interpretation model can predict the next sentence $S_{n+1}$ properly as illustrated in Figure 1 (right):
$\small S_{n+1}=f_{\theta}\Big{(}\sigma\big{(}W_{emb}(S_{n})+W_{pos}(S_{n})\big{)}\Big{)}=f_{\theta}(H_{real})\in\mathbb{Z}^{L}$ (6)
As a result, text synthesis is carried out as the seed passes through the seed interpretation model to predict the subsequent sentence.
Applying to Unconditional Text Synthesis
Figure 2: Illustration of text synthesizing method using the seed interpretation model in the inference phase.
Here, we assume that the training of the TESGAN framework, including adversarial training, is fully completed and describe how the seed interpretation model synthesizes text during the inference phase. Let $f_{\theta}^{*}(\cdot)$ and $g_{\phi}^{*}(\cdot)$ denote the frozen seed interpretation model and frozen generator, respectively. We can now synthesize text during the inference stage using the well-trained $f_{\theta}^{*}(\cdot)$ and $g_{\phi}^{*}(\cdot)$. At this point, $g_{\phi}^{*}(\cdot)$ will generate a fake seed $H_{fake}$ with the same dimensions as $H_{real}$, as shown in Equation 7. Then, $f_{\theta}^{*}(\cdot)$ can synthesize text using the fake seed, as shown in Figure 2. In other words, if the generator can skillfully create fake seeds $H_{fake}$ that imitate the distributions of $H_{real}$, then $H_{fake}$ can also generate appropriate subsequent sentences (a.k.a synthetic text). However, no matter how excellently $H_{fake}$ is generated by the generator, it is useless if it cannot be interpreted; therefore, training the seed interpretation model is crucial. We use the pre-trained GPT-2 Radford et al. (2018) [^2] https://huggingface.co/docs/transformers/model_doc/gpt2 model and fine-tune it with multi-turn text data to serve as the seed interpretation model. In addition, this model is only used to provide $H_{real}$ from the real text with frozen parameters during adversarial training. Thus, the seed interpretation model never affects the training of the generator and the discriminator during adversarial training, and vice versa. More detailed specifications of the seed interpretation model are explained in Appendix A.
Synthesizing text based on the generated fake seeds $H_{fake}$ by the generator is entirely different from autoregressive prompting. This is because prompting methods Wei et al. (2021); Ouyang et al. (2022); Chung et al. (2022) function by providing discrete tokens as input to a model that generates the next tokens based on the previous one. On the other hand, the generator of TESGAN creates continuous spaces $H_{fake}$ for synthesizing text from random noise, enabling unconditional text synthesis without explicit human instruction. Furthermore, research is actively being conducted to leverage continuous spaces (learnable query) for flexible model training Lester et al. (2021); Alayrac et al. (2022); Li et al. (2023); Dai et al. (2023). Most research explores various methods, including using a fixed learned query after model training or memorizing multiple learned queries and selecting them selectively as needed in different situations. However, the TESGAN framework differs from the mentioned studies in that its primary objective is to generate appropriate queries through the gernerator to produce appropriate sentences.